Please help me answer this question completely ( please don't answer if not complete), better by typing, if handwritten please write clearly. Thank you.

Subject: Inorganic Chemistry

Question: COUNT the number of valence electron per metal of the following compounds and then, ELUCIDE the structure.

Note:
bipy: Bipyridine
dppe: bis(diphenylphosphino)ethane

(a) Ru3 (H-CO)2 (CO)8 (bipy) (b) RuCl (PPH3) (n1-Ph2PCH2PPh2) (n5-C5H5)
Answer #1

Ans: * odd Rud) = Ru (dB) = 8 300 G - 2 Ru-Ree = 2 UCO) Valence shell = 10e electron Ru (2) = Ru (do) = 8 300 = 6 S 2 Ru-Ru 1